# Safe Vision-Language Models via Unsafe Weights Manipulation

### v1

@@ 3.2 SafeGround metrics
$\mathtt{P}^{t}=\left\{\begin{matrix}1&\text{if}\,f_{\texttt{VLM}}(v^{i}_{s},t^{i})>f_{\texttt{VLM}}(v^{i}_{u},t^{i})\\ 0&\text{otherwise}\\ \end{matrix}\right.$ (3)
@@
$\mathtt{P}^{v}=\left\{\begin{matrix}1&\text{if}\,f_{\texttt{VLM}}(v^{i},t^{i}_{s})>f_{\texttt{VLM}}(v^{i},t^{i}_{u})\\ 0&\text{otherwise}\\ \end{matrix}\right.$ (4)
@@ 3.3 Does fine-tuning improve safety?
Specifically, we prompt the original CLIP model and its safer counterpart Safe-CLIP [49] with unsafe and safe queries and report the introduced preference metrics $\mathtt{P}^{t}_{s}$, $\mathtt{P}^{t}_{u}$, $\mathtt{P}^{v}_{s}$, and $\mathtt{P}^{v}_{u}$ in Fig. 2. From the results, we can draw two main conclusions. On one hand, Safe-CLIP demonstrates to be significantly safer when tested on unsafe queries ($\mathtt{P}^{t}_{u}$, $\mathtt{P}^{v}_{u}$). However, surprisingly, Safe-CLIP degrades safety when tested on safe queries in both modalities, i.e., -$23\%$ in $\mathtt{P}^{t}_{s}$ and -$5.9\%$ in $\mathtt{P}^{v}_{s}$. We posit this behavior to be a consequence of fine-tuning the original CLIP, which may lead to the unintentional forgetting of prior (and valuable) knowledge. Note that these findings would have been harder to uncover using existing metrics as retrieval-based safety metrics [49] focus on both correctness and safety, as described in Sec. 3.1.
@@
Figure 3: Unsafe Weights Manipulation (UWM) discovers unsafe weights in the parameter space of a given VLM and manipulates them to improve model safety. First, UWM leverages safe and unsafe data, analyzing the model’s activations to estimate layer-wise safe and unsafe scores. These scores are then aggregated to produce a final score. Lastly, UWM targets and manipulates a small percentage of the top-scoring weights, making the original model safer.
@@ 4 Unsafe Weights Manipulation
An interesting outcome of Sec. 3.3 is that the original VLM has useful, safe behaviors that fine-tuning may inadvertently override. Here, we explore whether we can improve the safety of a model without training. Our method consists of three steps (Fig. 3): (i) estimating scores for safe and unsafe activations, (ii) comparing these scores to identify candidate weights, and (iii) modifying them toward safety.
@@
After collecting safe and unsafe scores, we aggregate them, quantifying the parameter’s influence on the model’s unsafe behavior as their ratio:
@@ 5.1 Implementation details
Hyperparameters. UWM introduces three main hyperparameters: which layers to prune, the cumulative sparsity score $\tau$, and the weight manipulation constant $\alpha$. We create the calibration set $\mathcal{D}$ by randomly sampling $400$ tuples per concept from ViSU’s training set. We observe consistent results using different calibration sets (e.g., ±$\pm 0.1$ in $\mathtt{GS}$). Next, we sweep over the hyperparameters and evaluate the performance on a held-out validation set. We observe that the layers requiring modification vary depending on the VLM architecture and pretraining strategy; however, both $\tau$ and $\alpha$ generalize well across experiments. Thus, we fix $\alpha=-1$ and $\tau=0.02$ for the experiments. We analyze these hyperparameters in the Supp. Mat. This procedure is also applied to all the baselines, selecting their best configuration.
[…]
@@ 5.2 Quantitative Results
The case of G-Unsafe highlights the need for novel metrics. According to retrieval-based metric $\mathcal{T}_{u}\text{-}\mathcal{V}_{u+s}$, this method reduces safety for unsafe textual queries. However, by looking at $\mathtt{P}_{u}^{t}$ (i.e., preference of safe images when prompted with unsafe text), its safety improves. This is a direct example of the discussion in Sec. 3.1. The drop in retrieval performance does not necessarily reflect lower safety but rather a degradation of the model’s knowledge, damaging its retrieval capabilities. This is further confirmed by the low zero-shot performance ($43.3\%$). Thus, our metrics successfully decouple safety from retrieval performance, enabling a more precise evaluation. Moreover, they expose the lower safety of G-Unsafe for safe queries ($\mathtt{P}_{s}^{t}$, $\mathtt{P}_{s}^{v}$).
On the contrary, G-Safe-CLIP improves safety across all four metrics while preserving more knowledge than the previous method, confirming the findings of the retrieval-based metrics: the method improves model safety.
@@
SafeGround Metrics. We conclude by discussing the performance according to the SafeGround metrics (last columns of Tab. 1) that combine the preferences to analyze safety across modalities and input type.
When analyzing the modalities ($\mathtt{Txt}_{s}$ and $\mathtt{Img}_{s}$), all methods struggle more with text queries, as the image modality score ($\mathtt{Img}_{s}$) is consistently lower than its textual counterpart ($\mathtt{Txt}_{s}$). Interestingly, their gap tends to increase as models get safer: CLIP shows a $1.7\%$ difference, while G-Safe-CLIP, UWM, and Safe-CLIP show increasing gaps of $2.7\%$, $8.3\%$, and $9.8\%$, respectively. This suggests that both training-free and training-based methods produce safer outputs for images compared to texts.
$\mathtt{PS}$ and $\mathtt{PU}$ (i.e., safety according to safe/unsafe input type), show consistent patterns for both baselines, with improvements for unsafe inputs ($\mathtt{PU}$) while lower performance on safe ones ($\mathtt{PS}$). Moreover, this latter case also applies to Safe-CLIP, with a decrease of $21.6\%$, further confirming its unsafe performance on safe inputs. In contrast, UWM improves safety across both safe and unsafe inputs. Finally, all methods improve the group score $\mathtt{GS}$.
@@ 5.3 Analysis & Discussion
$\Phi^{\mathtt{uns}}$ $\frac{\Phi^{\mathtt{uns}}}{\Phi^{\mathtt{sf}}}$ Adapt Zero-Shot ()($\uparrow$)( ↑ ) SafeGround Metrics ()($\uparrow$)( ↑ )
@@
[…]
@@
Scoring Function. In Tab. 3, we ablate the choice of the scoring function for computing the importance scores $\Phi$. We compare three variants: (i) using unsafe scores only $\Phi^{\mathtt{uns}}$, (ii) incorporating the aggregation score from Eq. 10, and (iii) applying the adaptive selection of Eq. 11.
Relying solely on $\Phi^{\mathtt{uns}}$ degrades zero-shot performance (-$15.5$), but already improves safety (+$3.4$ $\mathtt{GS}$). Introducing the ratio $\Phi^{\mathtt{uns}}/\Phi^{\mathtt{sf}}$ further enhances safety (+$11.8$ $\mathtt{GS}$) but severely impacts zero-shot performance (-$23.6$), making the model unsuitable for downstream tasks. The best trade-off is achieved with adaptive selection, improving safety (+$4.5$ $\mathtt{GS}$) while minimizing zero-shot degradation (-$7.8$). Additional ablations are provided in the Supp. Mat.
@@
[…]
@@ 6 Conclusion
This work investigates the safety challenges of Vision-Language Models (VLMs). We examine existing metrics [49] and find that they have limitations in assessing safety, as they rely on retrieval-based evaluations. Therefore, we complement these metrics by introducing SafeGround, a novel suite of metrics specifically designed for safety evaluation. SafeGround exposes a critical issue: training-based safety alignment techniques, such as Safe-CLIP [49], can compromise safe representations, leading to unsafe behaviors on safe queries. To address this, we propose UWM, a training-free method that identifies and manipulates specific unsafe weights within VLMs by analyzing how their information flow varies between safe and unsafe content. Our results show that, while Safe-CLIP [49] improves safety on unsafe queries, it fails on safe ones and degrades model knowledge. In contrast, UWM achieves a better trade-off between safety and knowledge preservation, marking a promising first step toward effective, training-free safety mechanisms for VLMs. Finally, we hope that SafeGround can serve as a valuable tool to expose unsafe behaviours in VLMs.
Limitations. While UWM is applicable to various architectures and VLMs (Tab. 4 and Tab. 5), it relies on weight localization and manipulation. The complexity of large-scale VLMs challenges the isolation of unsafe weights without affecting model capabilities, as individual parameters may encode overlapping knowledge. Indeed, both UWM and Safe-CLIP [49] exhibit limitations, highlighting that safety remains an open challenge requiring further research. Lastly, we mostly focus on contrastive-based VLMs, leaving the exploration of other architectures for future research.
@@
In this supplementary material, we add further ablation and analyses. Specifically, in Appendix A we i) provide a comprehensive ablation across all components of UWM and ii) expand baseline descriptions. Finally, in Appendix B we show further qualitative results and failure cases.
$\Phi^{\mathtt{uns}}$ $\frac{\Phi^{\mathtt{uns}}}{\Phi^{\mathtt{sf}}}$ $\Phi^{\mathtt{uns}}-\Phi^{\mathtt{sf}}$ Adapt Zero-Shot ()($\uparrow$)( ↑ ) SafeGround Metrics ()($\uparrow$)( ↑ )
@@
Text Encoder Image Encoder Zero-Shot ()($\uparrow$)( ↑ ) SafeGround Metrics ()($\uparrow$)( ↑ )
@@ A.1 Ablation
In the first case, adaptively selecting weights using only the unsafe scores $\Phi^{\mathtt{uns}}$ (third vs second row) improves the zero-shot performance ($\mathcal{V}_{s}\text{-}\mathcal{T}_{s}$ from $24.3\%$ to $30.6\%$), however, the overall safety scores decrease ($\mathtt{GS}$ from $4.6\%$ to $3.4\%$). Moreover, this method makes the model more unsafe when prompted with safe data ($\mathtt{PS}$, from the original $67.5\%$ to $64.3\%$). This behavior is consistent with the same method without adaptivity (second row, dropping from $67.5\%$ to $60.9\%$), showing that including the safe scores $\Phi^{\mathtt{sf}}$ in the scoring function is crucial to disentangle weights associated to unsafe concepts. For this reason, we explore an alternative score, replacing the merging function $\frac{\Phi^{\mathtt{uns}}}{\Phi^{\mathtt{sf}}}$ with the difference $\Phi^{\mathtt{uns}}-\Phi^{\mathtt{sf}}$. This change aims to reduce the unsafe scores $\Phi^{\mathtt{uns}}$ when high safe scores $\Phi^{\mathtt{sf}}$ are present, allowing for negative values. This strategy (fourth row) shows good zero-shot performance ($31\%$ in $\mathcal{V}_{s}\text{-}\mathcal{T}_{s}$), while being capable of improving safety (e.g., $+1.4\%$ in $\mathtt{GS}$). However, this method (i) degrades safety with safe inputs (i.e., $\mathtt{PS}$), and (ii) performs worse than UWM (e.g., $+1.4\%$ in $\mathtt{GS}$ vs $+3.3\%$ of UWM). Fianlly, the adaptive selection method (fifth row) is capable of improving zero-shot performance preservation ($38.7\%$ in $\mathcal{V}_{s}\text{-}\mathcal{T}_{s}$) in line with results of the main paper, however, the scoring function fails in improving safety, with decreases in $\mathtt{Img_{s}}$ and $\mathtt{GS}$.
Magnitude priors. Following existing works [60], we treat the weights’ magnitudes as priors and test our method by incorporating them into our scoring function, i.e. we multiply the final score $\Phi$ by $\left|W\right|$. We ablate this in Tab. 7 where we apply the priors to each modality independently. This experiment shows that our method is capable of improving safety while preserving zero-shot performance when no prior is applied (first row, e.g. $+4.4\%$ $\mathtt{Txt}_{\mathtt{s}}$). When the prior is applied to the text encoder (second row), UWM greatly improves safety without severe performance degradation (e.g., $+12.7\%$ $\mathtt{Txt}_{\mathtt{s}}$ and $+6.1\%$ $\mathtt{Img}_{\mathtt{s}}$). However, when the same procedure is applied to the vision encoder (last two rows), we observe great performance degradation. This shows that the image encoder weights do not serve as a reliable prior for identifying crucial weights. Thus, during our experiments, we use the magnitude prior only on the text encoder.
@@
UWM finds unsafe weights while preserving zero-shot performance when applied to the text encoder’s value and output projection layers (last two blocks). In the case of the value projection layer (third block), the method improves safety when applied in conjunction with Fc2, value, or output layers of the image encoder (e.g. Fc2 $+1.5\%$ in $\mathtt{GS}$). However, it greatly decreases the zero-shot performance ($\mathcal{V}_{s}\text{-}\mathcal{T}_{s}$), showing that the value layer of the text encoder is sensitive to pruning and important to the text encoder.
The best results are achieved when UWM is applied to the output projection layer of the text encoder (last block). We can observe better zero-shot performance preservation while improving safety. Specifically, the best configuration is achieved when pruning simultaneously with the Fc2 of the image encoder, where we improve model safety (e.g. $+3.3\%$ in $\mathtt{GS}$) while achieving high performance preservation. A similar performance is obtained by applying UWM to the output projection layer of the image encoder (last row). In this case, we have similar zero-shot performance preservation of the previous configuration, however, we achieve lower safety improvements (e.g. $-1.5\%$ in $\mathtt{PS}$). Thus, we prune the Fc2 of the image encoder and the output projection layer of the text encoder across all experiments.
@@
Ablation on $\alpha$. We keep $\tau=0.02$, i.e. our final configuration for $\tau$, and ablate $\alpha$ in the range from $-1$ to $1$ with intervals of $0.1$. We show the results in Fig. 6 and report the zero-shot performance $\mathcal{V}_{s}\text{-}\mathcal{T}_{s}$, the text $\mathtt{Txt}_{\mathtt{s}}$ and image $\mathtt{Img}_{\mathtt{s}}$ modality metrics and the group score $\mathtt{GS}$.
@@
Modality Ablation. We further ablate UWM by applying it individually to each modality in Tab. 9. This experiment aims to analyze the relative contributions of each modality to the overall model performance. We observe that UWM significantly improves the text encoder nearly achieving its best results. When applied to the image encoder, the improvements are lower with $+1.2\%$ in $\mathtt{Txt}_{\mathtt{s}}$ and $+1.0\%$ in $\mathtt{PS}$ and a slight decrease in $\mathtt{Img}_{\mathtt{s}}$ and $\mathtt{PU}$. This further shows the complexity of pruning meaningfully the image encoder. Nevertheless, when UWM is applied to both encoders, we obtain the best performance, showing the effectiveness of pruning both the image and text encoders simultaneously.

### v2

@@ 1 Introduction
00footnotetext: Corresponding author: moreno.dinca@unitn.it
@@ 3.2 SafeGround metrics
[…]
@@
[…]
@@ 3.3 Does fine-tuning improve safety?
Specifically, we prompt the original CLIP model and its safer counterpart Safe-CLIP [49] with unsafe and safe queries and report the introduced preference metrics $\mathtt{P}^{t}_{s}$, $\mathtt{P}^{t}_{u}$, $\mathtt{P}^{v}_{s}$, and $\mathtt{P}^{v}_{u}$ in Fig. 2. From the results, we can draw two main conclusions. Safe-CLIP demonstrates to be significantly safer when tested on unsafe queries ($\mathtt{P}^{t}_{u}$, $\mathtt{P}^{v}_{u}$). However, surprisingly, Safe-CLIP degrades safety when tested on safe queries in both modalities, i.e., -$23\%$ in $\mathtt{P}^{t}_{s}$ and -$5.9\%$ in $\mathtt{P}^{v}_{s}$. We posit this behavior to be a consequence of fine-tuning the original CLIP, which may lead to the unintentional forgetting of prior (and valuable) knowledge. Note that these findings would have been harder to uncover using existing metrics as retrieval-based safety metrics [49] focus on both correctness and safety, as described in Sec. 3.1.
@@
Figure 3: Unsafe Weights Manipulation (UWM) discovers unsafe weights of a given VLM and manipulates them to improve safety. Using safe and unsafe data within a single modality (e.g., the image modality shown in the center), UWM analyzes each encoder’s activations to estimate layer-wise safe and unsafe scores, which are then aggregated. Lastly, a small percentage of the top-scoring weights is targeted and manipulated towards safety. Applied independently to each encoder, UWM prevents cross-modal interference during scoring.
@@ 4 Unsafe Weights Manipulation
An interesting outcome of Sec. 3.3 is that the original VLM has useful, safe behaviors that fine-tuning may inadvertently override. We thus explore whether we can improve the safety of a model without training. Our method consists of three steps (Fig. 3): (i) estimating scores for safe and unsafe activations, (ii) comparing these scores to identify candidate weights, and (iii) modifying them toward safety. This workflow is applied independently to each encoder.
@@
After collecting safe and unsafe scores, we aggregate them, quantifying the parameter’s influence on the encoder’s unsafe behavior as their ratio:
@@ 5.1 Implementation details
Hyperparameters. UWM introduces three main hyperparameters: which layers to prune, the cumulative sparsity score $\tau$, and the weight manipulation constant $\alpha$. We create the calibration set $\mathcal{D}$ by randomly sampling $400$ tuples per concept from ViSU’s training set. We observe consistent results using different calibration sets (e.g., $\pm 0.1$ in $\mathtt{GS}$). Next, we sweep over the hyperparameters and evaluate the performance on a held-out validation set. We observe that the layers requiring modification vary depending on the VLM architecture and pretraining strategy; however, both $\tau$ and $\alpha$ generalize well across experiments. Thus, we fix $\alpha=-1$ and $\tau=0.02$ for the experiments. We analyze these hyperparameters in the Supp. Mat. This procedure is also applied to all the baselines, selecting their best configuration.
Method Zero-Shot $\left(\uparrow\right)$ Text & Image Retrieval $\left(\uparrow\right)$ Basic Preference Metrics $\left(\uparrow\right)$ SafeGround Metrics $\left(\uparrow\right)$
@@ 5.2 Quantitative Results
The case of G-Unsafe highlights the need for novel metrics. According to retrieval-based metric $\mathcal{T}_{u}\text{-}\mathcal{V}_{u+s}$, this method reduces safety for unsafe textual queries. However, by looking at $\mathtt{P}_{u}^{t}$ (i.e., preference of safe images when prompted with unsafe text), its safety improves. This is a direct example of the discussion in Sec. 3.1. The drop in retrieval performance does not necessarily reflect lower safety but rather a degradation of the model’s knowledge, damaging its retrieval capabilities. This is further confirmed by the low zero-shot performance ($43.3\%$). Thus, our metrics successfully decouple safety from retrieval performance, enabling a more precise evaluation. Moreover, they expose the lower safety of G-Unsafe for safe queries ($\mathtt{P}_{s}^{t}$, $\mathtt{P}_{s}^{v}$). In contrast, G-Safe-CLIP improves safety across all four metrics while preserving more knowledge, confirming the findings of the retrieval-based metrics: the method improves safety.
@@
SafeGround Metrics. We conclude by discussing the performance according to the SafeGround metrics (last columns of Tab. 1) that combine the preferences to analyze safety across modalities and input type. When analyzing the modalities ($\mathtt{Txt}_{s}$ and $\mathtt{Img}_{s}$), all methods struggle more with text queries, as the image modality score ($\mathtt{Img}_{s}$) is consistently lower than its textual counterpart ($\mathtt{Txt}_{s}$). Interestingly, their gap tends to increase as models get safer: CLIP shows a $1.7\%$ difference, while G-Safe-CLIP, UWM, and Safe-CLIP show increasing gaps of $2.7\%$, $8.3\%$, and $9.8\%$, respectively. This suggests that both training-free and training-based methods produce safer outputs for images compared to texts. $\mathtt{PS}$ and $\mathtt{PU}$ (i.e., safety according to safe/unsafe input type), show consistent patterns for both baselines, with improvements for unsafe inputs ($\mathtt{PU}$) while lower performance on safe ones ($\mathtt{PS}$). Moreover, this latter case also applies to Safe-CLIP, with a decrease of $21.6\%$, further confirming its unsafe performance on safe inputs. In contrast, UWM improves safety across both safe and unsafe inputs. Finally, all methods improve the group score $\mathtt{GS}$.
@@ 5.3 Analysis & Discussion
$\Phi^{\mathtt{uns}}$ $\frac{\Phi^{\mathtt{uns}}}{\Phi^{\mathtt{sf}}}$ Adapt Zero-Shot $(\uparrow)$ SafeGround Metrics $(\uparrow)$
@@
Model Basic Preference Metrics $\left(\uparrow\right)$ SafeGround Metrics $\left(\uparrow\right)$
@@
Scoring Function. In Tab. 3, we ablate the scoring function for computing $\Phi$. We compare three variants: (i) using unsafe scores only $\Phi^{\mathtt{uns}}$, (ii) incorporating the aggregation score from Eq. 10, and (iii) applying the adaptive selection of Eq. 11. Relying solely on $\Phi^{\mathtt{uns}}$ degrades zero-shot performance (-$15.5$), but already improves safety (+$3.4$ $\mathtt{GS}$). Introducing the ratio $\Phi^{\mathtt{uns}}/\Phi^{\mathtt{sf}}$ further enhances safety (+$11.8$ $\mathtt{GS}$) but severely impacts zero-shot performance (-$23.6$), making the model unsuitable for downstream tasks. The best trade-off is achieved with adaptive selection, improving safety (+$4.5$ $\mathtt{GS}$) while minimizing zero-shot degradation (-$7.8$). Additional ablations are provided in the Supp. Mat.
@@
Model Unsafety $\left(\downarrow\right)$ Knowledge Preservation $\left(\uparrow\right)$
@@ 6 Conclusion
This work investigates the safety challenges of Vision-Language Models (VLMs). We examine existing metrics [49] and find that they have limitations in assessing safety, as they rely on retrieval-based evaluations. Therefore, we complement them by introducing SafeGround, a novel suite of metrics specifically designed for safety evaluation. SafeGround exposes a critical issue: training-based safety alignment techniques, such as Safe-CLIP [49], can compromise safe representations, leading to unsafe behaviors on safe queries. We propose UWM, a training-free method that identifies and manipulates unsafe weights in VLMs by analyzing how the information flow varies between safe and unsafe content. UWM achieves a better trade-off between safety and knowledge preservation, marking a promising first step toward training-free safety techniques for VLMs. Finally, we hope SafeGround will serve as a valuable resource for the community, enabling researchers to uncover unsafe behaviors in VLMs.
Limitations. While UWM is applicable to various architectures and VLMs (Tab. 4 and Tab. 5), it relies on weight localization and manipulation. The complexity of large-scale VLMs challenges the isolation of unsafe weights without affecting model capabilities, as individual parameters may encode overlapping knowledge. As both UWM and Safe-CLIP [49] exhibit limitations, VLM safety remains an open challenge. Lastly, we focus mainly on contrastive-based VLMs, leaving other architectures for future research.
Acknowledgments. This work was supported by the EU Horizon project “ELIAS - European Lighthouse of AI for Sustainability” (No. 101120237), MUR PNRR project FAIR (PE00000013) funded by the NextGenerationEU, and the FIS project GUIDANCE (Debugging Computer Vision Models via Controlled Cross-modal Generation) (No. FIS2023-03251).
@@
In this supplementary material, we provide additional ablations and analysis. Specifically, in Appendix A we i) provide a comprehensive ablation across all components of UWM and ii) expand baseline descriptions. Finally, Appendix B shows further qualitative results and failure cases.
$\Phi^{\mathtt{uns}}$ $\frac{\Phi^{\mathtt{uns}}}{\Phi^{\mathtt{sf}}}$ $\Phi^{\mathtt{uns}}-\Phi^{\mathtt{sf}}$ Adapt Zero-Shot $(\uparrow)$ SafeGround Metrics $(\uparrow)$
@@
Text Encoder Image Encoder Zero-Shot $(\uparrow)$ SafeGround Metrics $(\uparrow)$
@@ A.1 Ablation
In the first case, adaptively selecting weights using only the unsafe scores $\Phi^{\mathtt{uns}}$ (third vs second row) improves the zero-shot performance ($\mathcal{V}_{s}\text{-}\mathcal{T}_{s}$ from $24.3\%$ to $30.6\%$), however, the overall safety scores decrease ($\mathtt{GS}$ from $4.6\%$ to $3.4\%$). Moreover, this method makes the model more unsafe when prompted with safe data ($\mathtt{PS}$, from the original $67.5\%$ to $64.3\%$). This behavior is consistent with the same method without adaptivity (second row, dropping from $67.5\%$ to $60.9\%$), showing that including the safe scores $\Phi^{\mathtt{sf}}$ in the scoring function is crucial to disentangle weights associated to unsafe concepts. For this reason, we explore an alternative score, replacing the merging function $\frac{\Phi^{\mathtt{uns}}}{\Phi^{\mathtt{sf}}}$ with the difference $\Phi^{\mathtt{uns}}-\Phi^{\mathtt{sf}}$. This change aims to reduce the unsafe scores $\Phi^{\mathtt{uns}}$ when high safe scores $\Phi^{\mathtt{sf}}$ are present, allowing for negative values. This strategy (fourth row) shows good zero-shot performance ($31\%$ in $\mathcal{V}_{s}\text{-}\mathcal{T}_{s}$), while being capable of improving safety (e.g., $1.4\%$ in $\mathtt{GS}$). However, this method (i) degrades safety with safe inputs (i.e., $\mathtt{PS}$), and (ii) performs worse than UWM (e.g., $1.4\%$ in $\mathtt{GS}$ vs $3.3\%$ of UWM). Fianlly, the adaptive selection method (fifth row) is capable of improving zero-shot performance preservation ($38.7\%$ in $\mathcal{V}_{s}\text{-}\mathcal{T}_{s}$) in line with results of the main paper, however, the scoring function fails in improving safety, with decreases in $\mathtt{Img_{s}}$ and $\mathtt{GS}$.
Magnitude priors. Following existing works [60], we treat the weights’ magnitudes as priors and test our method by incorporating them into our scoring function, i.e. we multiply the final score $\Phi$ by $\left|W\right|$. We ablate this in Tab. 7 where we apply the priors to each modality independently. This experiment shows that our method is capable of improving safety while preserving zero-shot performance when no prior is applied (first row, e.g. $4.4\%$ $\mathtt{Txt}_{\mathtt{s}}$). When the prior is applied to the text encoder (second row), UWM greatly improves safety without severe performance degradation (e.g., $12.7\%$ $\mathtt{Txt}_{\mathtt{s}}$ and $6.1\%$ $\mathtt{Img}_{\mathtt{s}}$). However, when the same procedure is applied to the vision encoder (last two rows), we observe great performance degradation. This shows that the image encoder weights do not serve as a reliable prior for identifying crucial weights. Thus, during our experiments, we use the magnitude prior only on the text encoder.
@@
UWM finds unsafe weights while preserving zero-shot performance when applied to the text encoder’s value and output projection layers (last two blocks). In the case of the value projection layer (third block), the method improves safety when applied in conjunction with Fc2, value, or output layers of the image encoder (e.g. Fc2 $1.5\%$ in $\mathtt{GS}$). However, it greatly decreases the zero-shot performance ($\mathcal{V}_{s}\text{-}\mathcal{T}_{s}$), showing that the value layer of the text encoder is sensitive to pruning and important to the text encoder.
The best results are achieved when UWM is applied to the output projection layer of the text encoder (last block). We can observe better zero-shot performance preservation while improving safety. Specifically, the best configuration is achieved when pruning simultaneously with the Fc2 of the image encoder, where we improve model safety (e.g. $3.3\%$ in $\mathtt{GS}$) while achieving high performance preservation. A similar performance is obtained by applying UWM to the output projection layer of the image encoder (last row). In this case, we have similar zero-shot performance preservation of the previous configuration, however, we achieve lower safety improvements (e.g. $1.5\%$ in $\mathtt{PS}$). Thus, we prune the Fc2 of the image encoder and the output projection layer of the text encoder across all experiments.
@@
Ablation on $\alpha$. We keep $\tau=0.02$, i.e. our final configuration for $\tau$, and ablate $\alpha$ in the range from 1-1 to $1$ with intervals of $0.1$. We show the results in Fig. 6 and report the zero-shot performance $\mathcal{V}_{s}\text{-}\mathcal{T}_{s}$, the text $\mathtt{Txt}_{\mathtt{s}}$ and image $\mathtt{Img}_{\mathtt{s}}$ modality metrics and the group score $\mathtt{GS}$.
@@
Modality Ablation. We further ablate UWM by applying it individually to each modality in Tab. 9. This experiment aims to analyze the relative contributions of each modality to the overall model performance. We observe that UWM significantly improves the text encoder nearly achieving its best results. When applied to the image encoder, the improvements are lower with $1.2\%$ in $\mathtt{Txt}_{\mathtt{s}}$ and $1.0\%$ in $\mathtt{PS}$ and a slight decrease in $\mathtt{Img}_{\mathtt{s}}$ and $\mathtt{PU}$. This further shows the complexity of pruning meaningfully the image encoder. Nevertheless, when UWM is applied to both encoders, we obtain the best performance, showing the effectiveness of pruning both the image and text encoders simultaneously.